What is the density of a cylindrical bar of radius 5.4 mm and length 10.8 cm that has a mass of 68 g? (in kg m−3 to 2.s.f)

Note:

The volume  of a cylinder is given by  where  is the radius of the cylinder and  is the length.

The density  is calculated using  where  is the mass of the object and  is the volume.


1
Expert's answer
2021-03-25T14:07:48-0400

Volume of the bar:


"V={\pi}r^2h=3.1416\times(5.4\cdot10^{-3}m)^2\times0.108m=9.9\cdot10^{-6}m^3"


"Density=\frac{mass}{volume}=\frac{0.068kg}{9.9\cdot10^{-6}m^3}=6.9\cdot10^3kg/m^3"


Answer: "6.9\cdot10^3kg/m^3"